# Towards Optimal Distributed Edge Coloring with Small Palettes

### v1

@@ 1.3 Related work
While the majority of the distributed community has previously focused on developing algorithms using at least $12\Delta-$ 1 colors [1, 4, 13, 14, 15, 19, 25, 20, 29, 34, 35], there has been some notable progress below the greedy threshold. We already discussed some of the early works, which showed how to compute a $(2\Delta-2)$-edge coloring deterministically [37, 38] or a $(1+\varepsilon)\Delta$-edge coloring randomized [12]. On the lower bound side, the authors of [9] established that coloring the edges of a graph with at most $22\Delta-$ 2 colors takes at least $\Omega(\log_{\Delta}n)$ rounds deterministically and $\Omega(\log_{\Delta}\log n)$ rounds randomized. In Table 1 we attempt to give a concise overview over the algorithms using less than $12\Delta-$ 1 colors.
Randomized edge coloring via the distributed Lovász Local Lemma
After the turn of the millennium, the attention towards this problem dwindled down for a while, with the next significant progress appearing not until 2015 from Elkin, Pettie & Su [13]. They improve upon [12] by providing a randomized $(1+\varepsilon)\Delta$-edge coloring algorithm that reduces the problem to the distributed Lovász Local Lemma and runs in $\mathcal{O}(\log^{\ast}\Delta+\frac{\log n}{\Delta^{1-o(1)}})$ rounds. In both algorithms it is clear that $\varepsilon$ need not be constant, but it is not obvious how small it can be made as a function of $\Delta$. Optimizing for palette size, the authors of [9] give a randomized $(\Delta+\widetilde{\Theta}(\sqrt{\Delta}))$-edge coloring algorithm that runs in $\mathrm{poly}(\log n)$ rounds. This almost reaches the natural threshold $\Delta+\Theta(\sqrt{\Delta})$ for randomized coloring algorithms, which is the smallest palette size for which there still is a constant probability of being able to color any edge given a random feasible coloring of its neighborhood. Further improvements to variants of the distributed Lovász Local Lemma [28, 11] subsequently reduced the randomized complexity of $(1+\varepsilon)\Delta$-edge coloring further down to $\mathrm{poly}(\log\log n)$.
Deterministic edge coloring
In [26] the authors leverage maximum matchings in bipartite graphs to get a $\mathcal{O}(\Delta^{9}\cdot\mathrm{poly}(\log n))$-time deterministic algorithm for $(3\Delta/2)$-edge coloring. Their technique involves iteratively removing graphs of maximum degree $3$, which reduce the maximum degree of the original graph by at least $2$. By coloring each of those extracted graphs with a distinct set of three colors they end up with a $(3\Delta/2)$-edge coloring of the entire graph. This result has undergone a series of improvements since then. First, Harris [30] presented improved algorithms for hypergraph maximal matchings, culminating in a reduced runtime of $\widetilde{\mathcal{O}}(\Delta^{4}\cdot\log^{6}n)$ rounds. Next, via a novel algorithm for the hypergraph sinkless orientation problem [8] the complexity was pushed down even further to just $\mathcal{O}(\Delta^{2}\cdot\log n)$ rounds. Furthermore, by leveraging a splitting technique from [24, 23] this algorithm can also be turned into an $(3/2+\varepsilon)\Delta$-edge coloring algorithm running in $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$ rounds. In particular, this implies a $\mathcal{O}(\log^{3}n)$-round algorithm for $(2\Delta-2)$-edge coloring in general graphs, which is still the state-of-the-art even for this larger palette size. Another technique to obtain deterministic edge coloring algorithms that use fewer colors is the method of conditional expectation, which can be used to obtain a $(\Delta+\mathrm{poly}(\log n))$-edge coloring algorithm running in $\Delta+\mathrm{poly}(\log n)$ rounds [21].
$\Delta$-vertex coloring
Building on the seminal work in [37], the authors of [22] decreased the upper bound for the randomized complexity of the $\Delta$-coloring problem down to $\mathcal{O}(\log\Delta)+2^{\mathcal{O}(\sqrt{\log\log n})}$ after over two decades of no progress on this problem. Further, a very recent result by Bourreau, Brandt and Nolin [7] shows that $\Delta$-coloring can be solved in $\mathcal{O}(\log^{4}\Delta+\log^{2}\Delta\cdot\log n\log^{\ast}n)$ deterministic time and $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ time randomized. This also automatically yields corresponding $(2\Delta-2)$-edge coloring algorithms. While the time complexity of the deterministic algorithm does not yield an improvement over the $(3/2+\varepsilon)\Delta$-edge coloring algorithm from [8], the randomized version is still the current state-of-the-art for $(2\Delta-2)$-edge coloring. Finally we note that for $\Delta=\omega(\log^{21}n)$ the randomized complexity even collapses to $\mathcal{O}(\log^{\ast}n)$ due to [16].
Edge coloring via Vizing chains
Vizing’s theorem [43] states that every simple undirected graph of maximum degree $\Delta$ can be edge-colored using just $\Delta+1$ colors. Since $\Delta$ is an easy lower bound on the number of colors necessary, the edge-chromatic number of any graph falls into one of just two classes. In the last years there has been a considerable amount of work dedicated to distributed versions of Vizing’s theorem. In [42], the authors give a randomized $\mathrm{poly}(\Delta,\log n)$ algorithm for $(\Delta+2)$-edge coloring as well as a deterministic $\mathrm{poly}(\Delta,\log n)$ algorithm that uses $\Delta+\mathcal{O}(\log n/\log\log n)$ colors. In [6] the authors give a deterministic $(\Delta+1)$-edge coloring algorithm in $\widetilde{\mathcal{O}}(\mathrm{poly}(\Delta)\cdot\log^{5}n)$ rounds of the $\mathsf{LOCAL}$ model, as well as a $\mathcal{O}(\mathrm{poly}(\Delta)\cdot\log^{2}n)$-round randomized algorithm. Optimizing for palette size, there is a recent deterministic $(\Delta+1)$-edge coloring algorithm [5, 6] in $\widetilde{\mathcal{O}}(\Delta^{84}\cdot\log^{5}n)$ rounds. While the exponent in $\Delta$ in the runtime is certainly not optimized yet, there is a fundamental limit to this approach. The authors of [9] have shown that this approach requires recoloring subgraphs of diameter $\Omega(\Delta\cdot\log n)$ in the worst case.
@@ 1.4 Organization of the paper
In Section 3 we present the main technical condition under which partial $(2\Delta-2)$-edge colorings can be extended. We start Section 4 by presenting the simpler version of our new algorithm, which provides the reduction to MIS. Later on, in Section 4.5, we explain which subroutines can be replaced their relaxed counterparts with faster algorithms in order to achieve a reduction to $(2\Delta-1)$-edge coloring. Finally, in Section 5 we explain which subroutines need to be replaced in order to obtain the faster randomized reduction. We use the appendix to give a short overview over all the subroutines from the existing literature used throughout our algorithms to be used as a point of reference.
@@ 2 Preliminaries
For any positive integer $k>0$ we define the power graph $G^{k}=(V,E^{k})$ by adding an edge for every pair of vertices $u,v\in V$ with $\mathrm{dist}(u,v)\leq k$ to $E^{k}$. For any positive integer $k>0$ and any vertex $v\in V$ we define the $r$-hop neighborhood of $v$ to be $N^{k}(v)=N_{G^{k}}(v)$. For a subset of edges $F\subseteq E$ we write $G-F$ to denote the graph obtained by removing all edges in $F$ from $G$. Further, for a subset $V^{\prime}\subseteq V$, we denote by $G[V^{\prime}]=(V^{\prime},E^{\prime})$ the subgraph of $G$ induced by $V^{\prime}$. The edge set $E^{\prime}$ is given by the subset of edges with both endpoints in $V$. Similarly, for a subset of edges $E^{\prime}\subseteq E$ we define the edge-induced subgraph $G[E^{\prime}]=(V^{\prime},E^{\prime})$, where $V^{\prime}$ is given by the subset of vertices that are incident to at least one edge in $E^{\prime}$. We define the degree $\mathrm{deg}(e)$ of an edge $e\in E$ as the number of edges sharing an endpoint with $e$. A hypergraph is a generalization of a graph in which each edge can contain more than two vertices. The rank of a hyperedge $e$ is the number of vertices contained in $e$. A proper $c$-edge coloring is a function $\varphi:E\to[c]$ such that for all adjacent edges $e,e^{\prime}\in E$ it holds that $\varphi(e)\neq\varphi(e^{\prime})$. In this paper we will work a lot with partial edge colorings that assign colors only to a subset of edges.
@@
Let $G=(V,E)$ be a graph, $F\subseteq E$ and $\varphi:E\setminus F\to[2\Delta-2]$ a partial edge coloring of $G$. Then we call $\varphi^{\star}$ an extension of $\varphi$ to $F$ if $\varphi^{\star}$ is a proper $(2\Delta-2)$-edge coloring of $G$ and $\varphi^{\star}(e)=\varphi(e)$ for all $e\in E\setminus F$.
We say a color $c$ is available for an edge $e\in F$, if no adjacent edge already has color $c$.
Definition 2.2 (edge-degree-choosable).
A subgraph $H\subseteq G$ is edge-degree-choosable, if for every assignment of lists =$\mathcal{L}=\{L(e):e\in E(H)\}$, such that $|L(e)|\geq\deg(e)$ for all $e\in E(H)$, there exists a proper edge-coloring of $G$ with colors from $\mathcal{L}$.
Note that any edge-degree choosable subgraph $H\subseteq G$ can be $(2\Delta-2)$-edge colored no matter how the edges outside of $G$ are colored.
@@
For two integers $\alpha,\beta\geq 1$ an $(\alpha,\beta)$-ruling set for a subset $W\subseteq V$ is a set of nodes $S\subseteq V$ such that the distance between any two nodes in $S$ is at least $\alpha$ and any node $w\in W$ has a distance of at most $\beta$ to the closest node in $S$. If $V=W$, we call $S$ an $(\alpha,\beta)$-ruling set.
Ruling sets generalize maximal independent sets, which are $(2,1)$-ruling sets. In our algorithm we will use ruling sets as a tool in order to partition the vertex sets into clusters with some specific properties. Most importantly, we will need that each cluster $C$ there is a central vertex $r\in C$ such that the whole $k$-hop neighborhood around $r$ is contained in $C$, for some parameter $k>0$. This ensures that each cluster has sufficiently many vertices, which will become important when we try to assign two exclusive matching edges to each cluster.
@@
In the third step of our algorithm, our goal is to assign two exclusive edges to each cluster, such that the set of all such edges forms a matching in $G$. Starting with a maximal matching, we construct an auxiliary hypergraph $H$ that lets us modify the matching in order to satisfy this property. The vertices of the hypergraph are the clusters in $G$ and we want to make sure that each cluster gets exclusive access to two hyperedges. This can be seen as an instance of the hypergraph sinkless orientation problem, which is a natural generalization of the fundamental sinkless orientation problem.
@@
If the minimum degree $\delta=\min_{v\in V}\mathrm{deg}(v)$ and the maximum rank $r=\max_{e\in E}\mathrm{rank}(e)$ of a hypergraph $H=(V,E)$ satisfy $\delta>r$, then this problem can be solved efficiently using an algorithm from [8].

### v2

@@ 1.3 Related work
Distributed graph coloring is one of the most extensively studied problems in the field of distributed computing. Many foundational results are covered in the excellent book by Barenboim and Elkin [8]. In this section, we survey both, results for greedy and non-greedy edge coloring, highlighting a range of deterministic and randomized algorithms.
Greedy distributed edge coloring
Deterministic algorithms
In a seminal paper from 1987, Linial showed that any deterministic $3$-coloring algorithm on the $n$-cycle takes at least $1/2\cdot\log^{\ast}n$ rounds in the $\mathsf{LOCAL}$ model [54]. Naor later extended this lower bound to apply to randomized algorithms as well [59]. On the upper bound side, Linial also presented an $\mathcal{O}(\log^{\ast}n)$-round $\mathcal{O}(\Delta^{2})$-coloring algorithm [54]. Combined with the classical one-color-per-round reduction this directly yields a $(2\Delta-1)$-edge coloring algorithm in $\mathcal{O}(\Delta^{2}+\log^{\ast}n)$ rounds. With some clever tricks, subsequent works [6, 9, 10, 60, 53, 33, 57] have improved on this simple color reduction scheme, culminating in the current state-of-the-art for $(2\Delta-1)$-edge coloring using $\mathcal{O}(\log^{12}\Delta+\log^{\ast}n)$ rounds [4]. Another line of research focuses on the complexity of graph coloring as a function of $n$, independent of $\Delta$. Early works by Panconesi & Srinivasan [61, 63] show that using network decompositions, one can find a $(2\Delta-1)$-edge coloring in $\exp(\mathcal{O}(\sqrt{\log n}))$ rounds. A long standing open question in the field asked for $(2\Delta-1)$-edge coloring algorithm in polylogarithmic time. Barenboim and Elkin made progress by showing that the $\Delta^{1+\varepsilon}$-coloring problem can be solved in $\mathcal{O}(\log^{2}n)$ rounds [7]. Ghaffari & Su later gave a $\mathrm{poly}(\log n)$ algorithm for $(2\Delta-1)(1+o(1))$-edge coloring [45], which in turn was improved by [42]. Fischer, Ghaffari, and Kuhn [31] fully resolved the question with a breakthrough $\mathcal{O}(\log^{9}n)$-round algorithm based on a novel hypergraph maximal matching procedure. Since then, the field has picked up the pace with a lot of new algorithms being developed in the last decade, e.g. [4, 29, 36, 37, 39, 43, 49, 50]. The fastest algorithm for general graphs currently comes from the MIS algorithm in [37] and runs in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds.
Randomized algorithms
The first randomized algorithm for greedy edge coloring follows from the seminal $\mathcal{O}(\log n)$-round MIS algorithm independently developed by Luby [56] and Alon, Babai and Itai [1]. This upper bound stood for almost three decades until it was surpassed by the $\mathcal{O}(\log\Delta)+2^{\mathcal{O}(\sqrt{\log\log n})}$-round algorithm introduced in [11], pioneering the now ubiquitous shattering framework for randomized algorithms. This algorithm was later improved to $\mathcal{O}(\log^{*}\Delta)+T^{\mathrm{det}}_{\mathrm{deg+1}}(\mathrm{poly}\log n)$ by Chang, Li, and Pettie [20]. Here, $T^{\mathrm{det}}_{\mathrm{deg+1}}(n^{\prime})$ is the deterministic complexity of the $(\mathrm{deg}+1)$-list coloring problem on instances of size $n^{\prime}$ [21]. A major breakthrough came with the first polylogarithmic network decomposition algorithm [65], reducing the complexity of the second term in both results and enabling $(2\Delta-1)$-edge coloring in $\mathrm{poly}(\log\log n)$ rounds. Subsequent works [29, 36, 37, 38, 64] continued to refine the runtime with the current state-of-the-art standing at $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds [37].
Non-greedy distributed edge coloring
@@
Despite the more challenging nature of the problem, there has also been a lot of progress in the non-greedy regime of coloring with $k<2\Delta-1$ colors. On the lower bound side, the authors of [17] established that coloring the edges of a graph with at most $22\Delta-$ 2 colors takes at least $\Omega(\log_{\Delta}n)$ rounds deterministically and $\Omega(\log_{\Delta}\log n)$ rounds randomized. While most of the algorithms breaking the greedy color barrier were only developed in the last decade, there are two remarkable results before the turn of the millennium.
$\Delta$-vertex coloring algorithms
In 1992, Panconesi & Srinivasan [61, 62] presented the first distributed $\Delta$-coloring algorithm running in $\exp(\mathcal{O}(\sqrt{\log n}))$ rounds. This deterministic result also implies the first distributed $(2\Delta-2)$-edge coloring algorithm. After two decades without improvements, the authors of [40] finally decreased the upper bound for the randomized complexity of $\Delta$-coloring to $\mathcal{O}(\log\Delta)+2^{\mathcal{O}(\sqrt{\log\log n})}$. Combining these seminal results with the new polylogarithmic algorithms for network decomposition [65, 38, 36, 29, 64, 37] yields runtimes of $\mathrm{poly}(\log n)$ and $\mathcal{O}(\log\Delta)+\mathrm{poly}(\log\log n)$, respectively. Recently, Bourreau, Brandt and Nolin [14] showed that $\Delta$-coloring can be solved in $\mathcal{O}(\log^{4}\Delta+\log^{2}\Delta\cdot\log n\log^{\ast}n)$ deterministic time and $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ randomized time. The randomized version of this result is also the current state-of-the-art for $(2\Delta-2)$-edge coloring. Moreover, for $\Delta=\omega(\log^{21}n)$ the randomized complexity of $\Delta$-coloring collapses to $\mathcal{O}(\log^{\ast}n)$ due to [32]. Finally, a very recent result [51] gives an optimal $\Delta$-coloring algorithm for a restricted graph class. However, since line graphs are not included in this class, it does not yield any new edge coloring results.
Randomized algorithms via the distributed Lovász Local Lemma
In 1998, Dubhashi, Grable & Panconesi [27] presented an $(1+\varepsilon)\Delta$-edge coloring algorithm using the Rödl nibble method, which runs in $\mathcal{O}(\log n)$ rounds, assuming $\Delta>(\log n)^{1+\gamma}$. After the turn of the millennium, there has been no progress on the problem, until Elkin, Pettie & Su [28] improved upon [27] in 2015. Their randomized $(1+\varepsilon)\Delta$-edge coloring algorithm reduces the problem to the distributed Lovász Local Lemma and runs in $\mathcal{O}(\log^{\ast}\Delta+\frac{\log n}{\Delta^{1-o(1)}})$ rounds. In both cases, $\varepsilon$ need not be constant, but it is not clear how small it can be made as a function of $\Delta$. Focusing on palette size over speed, the authors of [17] give a randomized $(\Delta+\widetilde{\Theta}(\sqrt{\Delta}))$-edge coloring algorithm that runs in $\mathrm{poly}(\log n)$ rounds—nearly matching the natural threshold $\Delta+\Theta(\sqrt{\Delta})$, the smallest palette size for which there still is a constant probability of being able to color any edge given a random feasible coloring of its neighborhood. Recently, new advances in variants of the distributed Lovász Local Lemma [48, 26] have reduced the complexity of $(1+\varepsilon)\Delta$-edge coloring further to $\mathrm{poly}(\log\log n)$.
Deterministic algorithms
In a 2018 paper [44], the authors present a $\mathcal{O}(\Delta^{9}\cdot\mathrm{poly}(\log n))$-time algorithm for $(3\Delta/2)$-edge coloring based on maximum matchings in bipartite graphs. Their result has since seen a series of improvements. First, Harris [50] presented improved algorithms for hypergraph maximal matchings, reducing the runtime to $\widetilde{\mathcal{O}}(\Delta^{4}\cdot\log^{6}n)$ rounds. Next, a novel algorithm for hypergraph sinkless orientation [15] further reduced the complexity to $\mathcal{O}(\Delta^{2}\cdot\log n)$ rounds. Using a splitting technique from [42, 41] this algorithm can also be adapted to a $(3/2+\varepsilon)\Delta$-edge coloring algorithm with runtime $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$. Notably, this implies a $\mathcal{O}(\log^{3}n)$-round algorithm for $(2\Delta-2)$-edge coloring in general graphs.
Vizing chain algorithms
A Vizing chain is an alternating path used to swap colors and resolve conflicts when trying to assign a new color to an edge in a graph that cannot be colored greedily. Vizing chains form a central role in Vizing’s celebrated result stating that every graph is $(\Delta+1)$-edge colorable [69]. Recent years have brought significant progress on distributed versions of Vizing’s theorem. In [68], the authors give the first $\mathrm{poly}(\Delta,\log n)$ algorithms for randomized $(\Delta+2)$-edge coloring and deterministic $\Delta+\mathcal{O}(\log n/\log\log n)$ edge coloring. Further, Bernshteyn and Dhawan [12, 13] recently gave a deterministic $(\Delta+1)$-edge coloring algorithm in $\widetilde{\mathcal{O}}(\Delta^{84}\cdot\log^{5}n)$ rounds, together with a faster randomized version running in $\mathcal{O}(\mathrm{poly}(\Delta)\cdot\log^{2}n)$ rounds. While the exponent in $\Delta$ in the runtime is certainly not yet optimized, there is a fundamental barrier to this approach. The authors of [17] have shown that it requires recoloring subgraphs of diameter $\Omega(\Delta\cdot\log n)$ in the worst case.
@@ 1.4 Organization of the paper
In Section 3 we present our main technical contribution: a sufficient condition under which partial $(2\Delta-2)$-edge colorings can be extended. Building on this, we introduce our new algorithms in Section 4. We start by presenting our reduction from $(2\Delta-2)$-edge coloring to MIS. Later on, in Section 4.6, we replace the MIS and maximal matching subroutines to achieve a reduction to just $(2\Delta-1)$-edge coloring. Then, in Section 5 we adapt the reduction to obtain a faster randomized algorithm. For convenience, we collect all subroutines from the literature that we use in our algorithms in the appendix, to serve as a point of reference.
@@ 2 Preliminaries
For any positive integer $k>0$ we define the power graph $G^{k}=(V,E^{k})$ by adding an edge for every pair of vertices $u,v\in V$ with $\mathrm{dist}(u,v)\leq k$ to $E^{k}$. For any positive integer $k>0$ and any vertex $v\in V$ we define the $k$-hop neighborhood of $v$ to be $N^{k}(v)=N_{G^{k}}(v)$. For a subset of edges $F\subseteq E$ we write $G-F$ to denote the graph obtained by removing all edges in $F$ from $G$. Further, for a subset $V^{\prime}\subseteq V$, we denote by $G[V^{\prime}]=(V^{\prime},E^{\prime})$ the subgraph of $G$ induced by $V^{\prime}$. The edge set $E^{\prime}$ is given by the subset of edges with both endpoints in $V^{\prime}$. Similarly, for a subset of edges $E^{\prime}\subseteq E$ we define the edge-induced subgraph $G[E^{\prime}]=(V^{\prime},E^{\prime})$, where $V^{\prime}$ is given by the subset of vertices that are incident to at least one edge in $E^{\prime}$. We define the degree $\mathrm{deg}(e)$ of an edge $e\in E$ as the number of edges adjacent to $e$. A hypergraph is a generalization of a graph in which each edge can contain more than two vertices. The rank of a hyperedge $e$ is the number of vertices contained in $e$. A proper $c$-edge coloring is a function $\varphi:E\to[c]$ that assigns different colors to adjacent edges. In this paper we will work a lot with partial edge colorings that assign colors only to a subset of edges.
@@
Let $G=(V,E)$ be a graph, $F\subseteq E$ and $\varphi:E\setminus F\to[2\Delta-2]$ a partial edge coloring of $G$. Then we call $\varphi^{\star}$ an extension of $\varphi$ to $F$ if $\varphi^{\star}$ is a proper $(2\Delta-2)$-edge coloring of $G$ and $\varphi^{\star}(e)=\varphi(e)$ for all $e\in E\setminus F$. Further, we say a color $c$ is available for an edge $e\in F$, if no adjacent edge already has color $c$.
@@
For two integers $\alpha,\beta\geq 1$, an $(\alpha,\beta)$-ruling set for a subset $W\subseteq V$ is a set of nodes $S\subseteq V$ such that the distance between any two nodes in $S$ is at least $\alpha$ and any node $w\in W$ has a distance of at most $\beta$ to the closest node in $S$. If $V=W$, we call $S$ an $(\alpha,\beta)$-ruling set.
Ruling sets generalize maximal independent sets, which are $(2,1)$-ruling sets. In our algorithm we will use ruling sets as a tool to partition the vertex sets into clusters with some specific properties. Most importantly, we will need that for each cluster $C$ there is a central vertex $r\in C$ such that the whole $k$-hop neighborhood around $r$ is contained in $C$, for some parameter $k>0$. This ensures that each cluster has sufficiently many vertices, which will become important when we try to assign two exclusive matching edges to each cluster.
@@
In the third step of our algorithm, our goal is to assign two exclusive edges to each cluster, such that the set of all such edges forms a matching in $G$. Starting with a maximal matching, we construct an auxiliary hypergraph $H$ that lets us modify the matching in order to satisfy this property. This can be seen as an instance of the hypergraph sinkless orientation problem, which is a natural generalization of the fundamental sinkless orientation problem.
@@
If the minimum degree $\delta=\min_{v\in V}\mathrm{deg}(v)$ and the maximum rank $r=\max_{e\in E}\mathrm{rank}(e)$ of $H$ satisfy $\delta>r$, then this problem can be solved efficiently using the algorithms in Theorems A.8 and A.9. We follow the standard assumption that each vertex in the $\mathsf{LOCAL}$ model knows the number of nodes $n=\lvert V\rvert$ of the graph, as well as the maximum degree $\Delta$.